10. Chapter 10, Problem 10 Aa Aa In 1974, Elizabeth Loftus and John Palmer conducted a classic study demonstrating how the language used to ask a question can influence eyewitness memory. In the study, college students watched a film of an automobile accident and then were asked questions about what they saw. One group was asked, "About how fast were the cars going when they smashed into each other?" Another group was asked the same question except the verb was changed to "hit" instead of "smashed into." The "smashed into" group reported significantly higher estimates of speed than the "hit" group. [Loftus, G. R. & Palmer, J. C. (1974). Reconstruction of automobile destruction: An example of the interaction between language and memory. Journal of Verbal Learning & Verbal Behavior, 13, 585-589.] Suppose a researcher repeats this study with a sample of today's college students and obtains the following results: Estimated Speed "Hit" "Smashed Into" n 15 n = 15 M 40.8 M 34.0 SS = 414 SS 510 Do the results indicate a significantly higher estimated speed for the "smashed into" group? Use a one-tailed test with a .01. (Reminder: Choose the correct sign for the critical value, and use three decimal places for the critical values, two for the estimated standard error and t statistic.) t Distribution Degrees of Freedom 21 -1.0 3.0 -2.0 0.0 1.0 2.0 3.0 X Pooled Variance Estimated Standard Error t Statistic Critical Values
Conclusion: O Fail to reject Ho. Participants who hear "smashed into" rather than "hit" estimate significantly higher speeds O Fail to reject Ho. Participants who hear "smashed into" rather than "hit" do not estimate significantly higher speeds O Reject Ho. Participants who hear "smashed into" rather than "hit" estimate significantly higher speeds. O Reject Ho. Participants who hear "smashed into" rather than "hit" do not estimate significantly higher speeds. Compute the estimated value for Cohen's d to measure the size of the effect. (Use two decimal places.) Estimated d = Select the correct item from each pulldown menu to complete a sentence demonstrating how the results of the hypothesis test and the measure of effect size would appear in a research report. The results that participants who heard the verb "smashed into" estimated significantly higher speeds than those who heard the verb "hit,"
Answer #1

For Smashed Into :

Sample mean, x̅1 = 40.8

Sample standard deviation, s1 = SQRT(SS/(n-1) = 6.0356

Sample size, n1 = 15

For Hit :

Sample mean, x̅2 = 34

Sample standard deviation, s2 = SQRT(SS/(n-1) = 5.4380

Sample size, n2 = 15

α = 0.01

Null and Alternative hypothesis:

Ho : µ1 = µ2

H1 : µ1 > µ2

Pooled variance :

S²p = ((n1-1)*s1² + (n2-1)*s2² )/(n1+n2-2) = ((15-1)*6.0356² + (15-1)*5.438²) / (15+15-2) = 33

Standard error, se = √(s²p(1/n1 + 1/n2 ) = √(33*(1/15 + 1/15)) = 2.0976

Test statistic:

t = (x̅1 - x̅2) / se = (40.8 - 34) / 2.0976 = 3.24

df = n1+n2-2 = 28

Critical value :

Right tailed critical value, t crit = ABS(T.INV(0.01, 28)) = 2.467

Decision:

t = 3.24 >2.467, Reject the null hypothesis

Conclusion:

Reject Ho. Participants who hear "smashed into" rather than "hit" estimate significantly higher speeds.

Cohen's d = (x̅1 - x̅2)/√s²p = (40.8-34)/√33 = 1.18